### v2

# Analyzing Fairness of Computer Vision and Natural Language Processing Models

Abstract
Machine learning (ML) algorithms play a crucial role in decision-making across diverse fields such as healthcare, finance, education, and law enforcement. Despite their widespread adoption, these systems raise ethical and social concerns due to potential biases and fairness issues. This study focuses on evaluating and improving the fairness of Computer Vision and Natural Language Processing (NLP) models applied to unstructured datasets, emphasizing how biased predictions can reinforce existing systemic inequalities. A publicly available dataset from Kaggle was utilized to simulate a practical scenario for examining fairness in ML workflows.
To address and mitigate biases, the study employed two leading fairness libraries: Fairlearn by Microsoft, and AIF360 by IBM. These tools offer comprehensive frameworks for fairness analysis, including metrics evaluation, result visualization, and bias mitigation techniques. The research aims to measure bias levels in ML models, compare the effectiveness of these fairness libraries, and provide actionable recommendations for practitioners.
The results demonstrate that each library possesses distinct strengths and limitations in evaluating and mitigating fairness. By systematically analyzing these tools, the study contributes valuable insights to the growing field of ML fairness, offering practical guidance for integrating fairness solutions into real-world applications. This research underscores the importance of building more equitable and responsible machine learning systems.
1 Introduction
Machine learning algorithms are extensively employed across various fields, including entertainment, shopping, healthcare, finance, education, law enforcement, and critical areas such as loan approvals [1] and hiring decisions [2, 3]. They offer benefits like consistent performance and the ability to analyze numerous variables [4, 5]. However, these systems are not immune to biases, which can lead to unfair outcomes [6, 7]. Biases in machine learning are particularly concerning when decisions directly affect individuals or communities, potentially leading to discriminatory results. Ensuring that these systems operate ethically and equitably is crucial. Fair decision-making necessitates impartiality, avoiding favoritism based on inherent or acquired characteristics. Biased algorithms deviate from this principle, skewing decisions in favor of certain groups.
The idea of ”fairness” in algorithmic systems is deeply tied to the sociotechnical context. Different forms of fairness-related harms have been identified:
1.
Allocation Harm: Unequal distribution of opportunities or resources, such as an algorithm disproportionately favoring men over women for job offers [8].
2.
Quality-of-Service Harm: Unequal performance across groups, e.g., facial recognition systems misclassifying Black women more frequently than White men [9], or speech recognition systems underperforming for users with speech disabilities [10].
3.
Stereotyping Harm: Reinforcing societal stereotypes, such as image searches for ”CEO” showing predominantly White men [8].
4.
Denigration Harm: Producing offensive outputs, such as misclassifying individuals as animals or chatbots generating slurs [8].
5.
Representation Harm: Over- or under-representing certain groups, such as racial bias in welfare fraud investigations or neglecting elderly populations in surveillance applications [8].
6.
Procedural Harm: Violations of social norms in decision-making practices, such as penalizing job candidates for extensive experience or failing to ensure transparency and accountability in algorithmic decisions [11].
These harms often overlap and highlight the importance of addressing fairness concerns from the development phase of machine learning systems.
This study examines the fairness of machine learning models applied to structured datasets in classification tasks. Structured datasets, known for their organized format, facilitate efficient analysis and are frequently used in ML applications. However, these datasets often reflect historical biases or systemic discrimination, which can influence the fairness of machine learning models, making robust evaluation and mitigation essential.
For this research, a publicly available dataset [12] from Kaggle was chosen. Kaggle datasets offer realistic scenarios for evaluating machine learning models and are well-suited for fairness studies. The dataset underwent preprocessing and was used to develop classification models, which involve predicting discrete labels based on input features. Classification tasks are particularly significant in fairness research since biased predictions can disproportionately impact specific groups.
To identify and address biases, the study employed three advanced fairness libraries: Fairlearn by Microsoft, AIF360 by IBM, and the What-If Tool by Google. These libraries provide tools for assessing fairness metrics, visualizing model behavior, and applying bias mitigation techniques. Leveraging these tools, the research systematically evaluates the fairness of classification models and investigates strategies to minimize bias.
The objectives of this study are:
1.
To measure the level of bias present in machine learning models trained on the chosen dataset.
2.
To compare the performance and effectiveness of the three fairness libraries in identifying and mitigating bias.
3.
To offer actionable insights on implementing fairness tools in real-world machine learning workflows.
The remainder of this paper is structured as follows: Section 2 reviews related work on machine learning fairness. Section 3 details the methodology, including dataset preprocessing and model development. Section 4 discusses fairness analyses using the selected libraries and their functionalities. Section 5 presents a comparative evaluation and results of the libraries. Finally, Section 6 concludes with a summary of the findings.
2 Review of Related Work
Bias in machine learning (ML) models is an increasingly pressing concern, particularly as these models influence critical societal sectors like healthcare, hiring, and criminal justice. A growing body of research has examined the origins, manifestations, and mitigation strategies of bias, offering a solid foundation for tackling this pervasive issue.
One prominent area of study focuses on identifying and categorizing biases in ML models. Ref. [13] outlines a taxonomy of biases, classifying them into historical, representation, and measurement biases. Historical bias arises from pre-existing inequities in the data, even before applying ML methods. Representation bias occurs when certain groups are under- or over-represented in the training data, leading to skewed predictions [14]. Measurement bias results from inaccuracies in the features or labels used for training due to flawed measurement processes.
Another line of research addresses methods for detecting bias. Tools such as disparate impact analysis [15] and fairness metrics like demographic parity, equal opportunity, and disparate mistreatment [16] are widely used. In structured datasets, researchers focus on two fairness dimensions: group fairness, which ensures equitable treatment across demographic groups, and individual fairness, which emphasizes treating similar individuals similarly [17]. Ref. [18] explores the conflicts between these fairness definitions, highlighting the need for trade-offs tailored to specific contexts.
The technical challenges of mitigating bias have also been extensively studied. Pre-processing techniques, such as re-weighting data samples or modifying labels to enhance fairness, address biases at the dataset level [19]. In-processing methods, such as adversarial debiasing [20], embed fairness constraints directly into model training. Post-processing approaches modify model outputs to meet fairness criteria, including re-ranking methods proposed by [16]. However, Ref. [21] illustrates the inherent trade-offs between fairness and accuracy, demonstrating the difficulty of optimizing both simultaneously.
Bias analysis in structured datasets has received significant attention due to the prevalence of tabular data in decision-making systems. Such datasets often carry latent biases reflecting historical inequities or systemic discrimination. For instance, the COMPAS dataset, used in criminal justice, exhibits racial disparities in predictive outcomes [22]. Research has also highlighted how feature selection and preprocessing can either mitigate or exacerbate biases. Ref. [23] discusses the impact of feature correlation with sensitive attributes on fairness and proposes methods to disentangle these relationships.
Recent advances emphasize interpretability as a critical component of bias analysis. Ref. [24] introduced LIME (Local Interpretable Model-agnostic Explanations), enabling stakeholders to understand and detect biased patterns in model predictions. Similarly, Ref. [25] developed SHAP (SHapley Additive exPlanations), providing consistent and accurate feature importance values. These interpretability tools have been instrumental in identifying biases in structured datasets by allowing detailed analysis of how features contribute to unfair outcomes.
Intersectionality has also emerged as a key consideration in bias studies. Ref. [26] highlights the necessity of evaluating models across multiple demographic dimensions, revealing compounded disparities for intersectional groups, such as Black women in facial recognition systems. For structured datasets, Ref. [27] proposes fairness-enhancing strategies that account for multiple subgroups simultaneously, avoiding the limitations of single-axis fairness analyses.
The literature on bias in machine learning encompasses a broad spectrum of topics, from foundational concepts and detection methods to mitigation strategies and interpretability tools. Despite significant progress, challenges remain in applying these techniques to structured datasets, particularly in balancing fairness with competing goals like accuracy and interpretability. This review emphasizes the need for continued research into holistic, context-sensitive approaches for understanding and addressing bias in ML models.
3 Dataset and Model Details
For the computer vision model, we used the UTKFace dataset [12]. The UTKFace dataset is a collection of over 20,000 facial images labeled with demographic information, including age, gender, and ethnicity. It is widely used for machine learning tasks such as facial analysis and demographic prediction. The dataset includes images cropped and aligned to facilitate consistent analysis across models.
In terms of machine learning applications, this dataset is commonly utilized for tasks like age and gender prediction, facial recognition, and demographic analysis. Researchers often use deep learning techniques, including convolutional neural networks (CNNs), to extract features and train models on this dataset for predictive tasks.
For the Natural Language Processing (NLP) model, we used the California Independent Medical Review (IMR) Dataset [12]. The California Independent Medical Review (IMR) Dataset available on Kaggle contains healthcare-related data focusing on medical review decisions. It is primarily used for Natural Language Processing (NLP) tasks and contains text-based reviews related to insurance appeals. The dataset is valuable for modeling tasks that involve text classification, sentiment analysis, or decision-making predictions in a healthcare context.
While the dataset itself consists of text entries, researchers and developers may employ machine learning models such as Logistic Regression, Random Forest, or advanced NLP models like BERT for tasks such as understanding decision outcomes or identifying patterns in appeals based on textual data.
4 Implementation of Fairness Analyses
In recent years, the growing use of machine learning (ML) across various sectors has raised significant concerns regarding fairness and bias in classification models. Given their potential impact on decision-making in critical fields such as healthcare, finance, and criminal justice [28], ensuring fairness in these models has become crucial. Bias in ML models can arise from factors like imbalanced training data, model selection, and inherent societal biases, ultimately resulting in discriminatory outcomes against marginalized groups [29].
To address these challenges, several libraries and tools have been created to help practitioners analyze and reduce bias in their models. Notable among them are Fairlearn, AIF360, and the What-If Tool, each offering unique features for evaluating and enhancing fairness.
Fairlearn [30]: Developed by Microsoft, this toolkit assists data scientists in assessing and improving the fairness of AI models. It provides a range of metrics to evaluate fairness and algorithms to mitigate unfairness. Fairlearn emphasizes both the social and technical dimensions of fairness, helping users understand how various aspects of a model contribute to disparities among groups.
AIF360 [31]: Developed by IBM, AIF360 is a comprehensive library that offers a variety of metrics for fairness assessment and techniques for mitigating bias across the entire AI application lifecycle. It supports integration at various stages of the machine learning pipeline, promoting fairness-aware modeling. The library includes metrics for evaluating fairness across societal demographics and offers re-parameterization strategies to improve model robustness.
What-If Tool [32]: Created by Google, this interactive visualization tool enables users to explore and analyze ML models without needing to write any code. It facilitates testing model performance in hypothetical scenarios, helping users understand and explain model behavior. The tool allows users to examine how different demographics influence model predictions, providing a deeper understanding of potential bias.
These tools are vital for researchers aiming to identify and mitigate bias in ML classification models, offering the methodologies necessary to ensure equitable AI systems and informed decision-making processes.
5 Results and Discussions
This section presents the results and outcomes of applying fairness libraries to computer vision (CV) and natural language processing (NLP) models. The primary goal of this analysis is to improve fairness metrics while maintaining or enhancing performance metrics. Detailed code and implementation specifics are provided in [33].
For the CV model, the sensitive features analyzed for potential bias are gender and ethnicity, with the target variable being the prediction of a person’s age group. For the NLP model, the target variable is the determination of whether a patient requires urgent intervention, based on a combination of the doctor’s findings, patient gender, and age group. The sensitive features in this case are gender and age group, with the aim of identifying and addressing any model bias related to these attributes.
The subsequent discussion highlights the results obtained from applying fairness libraries to both the CV and NLP models. Table 1 summarizes the strengths and weaknesses of the libraries we use in this analysis (Fairlearn and AIF360) in addition to another library called What-If-Tool created by Google which is not powerful enough to be used in this study.
Table 1: Pros and Cons of the three fairness libraries.
For Fairlearn, accuracy was used to evaluate the overall predictive performance of the income classification model, while demographic parity difference served as a fairness metric to measure disparities in outcomes across demographic groups. Mitigation strategies were implemented at three stages of the machine learning pipeline: preprocessing, in-processing, and postprocessing.
Preprocessing: Techniques like the Correlation Remover addressed biases by eliminating correlations between sensitive and non-sensitive features while preserving data integrity [34].
In-processing: Algorithms such as Exponentiated Gradient incorporated fairness constraints during model training to ensure fairer decision boundaries [35].
Postprocessing: Methods like the Threshold Optimizer adjusted predictions after model training to satisfy fairness criteria precisely, ensuring no residual disparities [36].
The evaluation identified the best algorithms as those balancing high predictive accuracy with reduced bias, demonstrating the need for an integrated fairness approach to mitigate biases in structured datasets.
For AIF360, mitigation algorithms were applied at multiple stages to address bias comprehensively:
Reweighing: It assigns weights to training instances to correct imbalances in demographic representation, addressing biases embedded in the dataset.
Equalized Odds: It imposes constraints during model training to align true positive and false positive rates across sensitive groups, enhancing fairness without significantly reducing accuracy.
Disparate Impact Ratio: Ratio of selection rates which is the percentage of samples with positive selection.
Learning Fair Representations (Consistency Score): The metric computes the consistency score. Individual fairness metric that measures how similar the (predicted) labels are for similar instances (records). It compares a model’s classification prediction of a given data item x to its k-nearest neighbors, kNN(x). It applies the kNN function to the full set of examples to obtain the most accurate estimate of each point’s nearest neighbors
Adversarial Debiasing: Adversarial debiasing is an in-processing technique that utilizes adversarial learning to mitigate bias. The method involves training a predictive model (the classifier) while simultaneously training an adversarial model that seeks to predict the sensitive attribute (e.g., gender or race) from the classifier’s predictions. The classifier is then optimized to minimize its primary objective (e.g., accuracy) while also minimizing the adversary’s ability to detect the sensitive attribute. This process reduces the dependency of the classifier’s predictions on sensitive features, thereby enhancing fairness.
Equalized Odds: Equalized odds is a fairness constraint that ensures the predictive performance of a model is consistent across all groups defined by a sensitive attribute. Specifically, it requires that the true positive rate (TPR) and false positive rate (FPR) are equal for all groups. By enforcing parity in these rates, equalized odds minimizes discrimination and ensures that the model’s predictive behavior is fair across different demographic groups.
Reject Option Classification (ROC): Reject option classification is a postprocessing technique that adjusts the decision boundary of a trained classifier. It introduces a ”reject option” for samples with a high likelihood of misclassification or bias. In this option, predictions for individuals from an unprivileged group with borderline scores are reclassified to ensure fairness. ROC aims to correct unfair outcomes by altering decisions in favor of fairness, often at the cost of slight performance reductions.
5.1 Bias Mitigation in Computer Vision Models: A Comparative Analysis of Fairlearn and AIF360
5.1.1 Fairlearn Library
The Fairlearn library was employed to evaluate fairness in a computer vision model designed for age group classification. The dataset consisted of four age groups: 0–18, 18–30, 30–80, and above 80 years. Since Fairlearn does not support multi-class classification directly, metrics were computed separately for each age group. Results are summarized in Figure 1.
1.
Metrics:
Demographic Parity Difference (DPD): This measures the difference in selection rates among groups within a sensitive feature, such as ethnicity. A small DPD indicates reduced bias, implying the model treats all groups equally.
Equalized Odds Difference (EOD): This evaluates the quality-of-service harm by comparing the model’s performance (e.g., false positive rates and true positive rates) across groups within a sensitive feature.
2.
Results:
For the age group 30–80, a DPD of 55% was observed, indicating significant bias. In contrast, bias was minimal for groups above 80 years.
Mitigation algorithms were applied to reduce bias:
Correlation Remover: Reduced bias moderately with a minimal drop in accuracy from 66% to 63%.
Exponentiated Gradient: Further reduced bias but decreased accuracy more significantly.
Threshold Optimizer: Eliminated bias almost entirely but reduced accuracy to 18%, indicating a severe trade-off between fairness and performance.
Grid Search: Demonstrated minimal impact on bias and decreased accuracy to 45%.
3.
Insights:
Reducing bias often results in reduced model accuracy.
The Correlation Remover and Exponentiated Gradient algorithms provided a balance between bias reduction and performance retention, making them preferable for this study.
Figure 1: The results of applying Fairlearn to the computer vision model.
5.1.2 AIF360 Library
The AIF360 library was used to analyze bias in the same computer vision problem, leveraging its capability to handle multi-class classification directly. Results are summarized in Figure 2. Five fairness metrics were analyzed:
1.
Metrics:
Statistical Parity Difference (SPD): Analogous to DPD, measures selection rate differences among sensitive groups.
Average Odds Difference (AOD): Averages the differences in false positive and true positive rates between groups.
Equal Opportunity Difference (EOD): Measures recall disparities between privileged and unprivileged groups.
Generalized Entropy Index: Quantifies prediction randomness across groups to ensure fairness.
Theil Index: Similar to the entropy index but more sensitive to small datasets.
2.
Results:
The accuracy of the original model was 70%, with significant bias observed across metrics.
Mitigation algorithms were applied:
Reweighing and Disparate Impact Remover: Reduced bias effectively while maintaining the original accuracy.
Adversarial Debiasing: Also showed promising results with minimal accuracy degradation.
Post-Processing Algorithms (Equalized Odds Postprocessing and Reject Option Classification): Performed poorly, significantly increasing some fairness metrics and reducing accuracy to 17–18%.
3.
Insights:
AIF360 provides a robust framework for bias analysis with multiple metrics and mitigation algorithms covering different stages of the machine learning lifecycle.
Reweighing and Disparate Impact Remover were the most effective algorithms, offering a balance between fairness and accuracy.
Technical Challenges
Some Fairlearn mitigation algorithms, such as Exponentiated Gradient Reduction, were incompatible with TensorFlow-based models.
The Learning Fair Representation algorithm required excessive computational resources, limiting its utility on local machines.
This comparative analysis underscores the trade-offs between bias reduction and model accuracy in fairness-focused machine learning. While Fairlearn offers simplicity for binary tasks, AIF360 provides a more comprehensive suite for multi-class classification. The study highlights the importance of selecting appropriate mitigation algorithms based on the specific requirements and constraints of the application.
Figure 2: The results of applying AIF360 to the computer vision model.
5.2 Bias Mitigation in NLP Models: Evaluation Using Fairlearn and AIF360
5.2.1 Fairlearn Library
The Fairlearn library was utilized to analyze fairness in a binary classification task within a natural language processing (NLP) model. Two fairness metrics were evaluated: Demographic Parity Difference (DPD) and Equalized Odds Difference (EOD). Results are summarized in Figure 3.
1.
Results:
The original model showed a high performance with an accuracy of 97–98%.
Grid Search: Increased bias, making it the least effective mitigation algorithm.
Correlation Remover and Exponentiated Gradient: Showed minimal changes to the fairness metrics compared to the original model, indicating limited efficacy.
Threshold Optimizer: Demonstrated the best performance among the mitigation algorithms. It reduced the EOD while maintaining the original accuracy, though it did not affect the DPD.
2.
Insights:
The Threshold Optimizer emerged as the most effective mitigation algorithm for reducing EOD without compromising performance.
Correlation Remover and Exponentiated Gradient had negligible impact on bias, while Grid Search was counterproductive, increasing bias metrics.
Figure 3: The results of applying Fairlearn to the NLP model.
5.2.2 AIF360 Library
The AIF360 library was similarly applied to assess fairness in the same NLP model, with the same metrics used for evaluation.
1.
Results:
The original model maintained an accuracy of 98%.
No observable changes in fairness metrics were noted after applying various mitigation algorithms.
The mitigation algorithms did not significantly influence either the fairness or performance metrics.
2.
Insights:
The lack of improvement in fairness metrics suggests that the mitigation algorithms in AIF360 may not be impactful for this specific NLP model.
Alternative models or configurations may yield better results with the AIF360 library.
In the context of NLP models, the Threshold Optimizer from Fairlearn showed some potential in reducing bias while preserving accuracy, particularly for EOD. However, the overall impact of mitigation algorithms in both Fairlearn and AIF360 was limited, highlighting the need for further exploration of model-specific or task-specific fairness solutions.
Figure 4: The results of applying AIF360 to the NLP model.
6 Conclusion
The study highlighted the role of biased predictions in reinforcing systemic inequalities and underscores the need for evaluating and improving fairness in ML workflows. By utilizing a Kaggle dataset, this research provided practical insights into how fairness libraries, such as Fairlearn by Microsoft and AIF360 by IBM, can be employed to assess, visualize, and mitigate bias in ML models.
The findings indicate that each fairness tool has its own strengths and limitations in addressing bias, offering valuable guidance for researchers and practitioners. These insights contribute to the growing field of ML fairness, reinforcing the necessity of integrating fairness solutions into real-world ML applications to build more equitable and responsible systems. The study ultimately underscores the need for continuous efforts in ensuring that ML models operate ethically and do not perpetuate existing societal inequalities.
Acknowledgement
We would like to thank Yale Center for Research Computing for supporting A.K. through the CAREERS project administered by the PSU Institute for Computational and Data Sciences (ICDS) under the National Science Foundation Award with No. 2018873.